OpenStudy (minthia):

Insert <,> or = to make the sentence true - 1/3 - 2/5

OpenStudy (danjs):

you can always put both dractions over the same denominator and simply compare the numerators...

OpenStudy (minthia):

huh?

OpenStudy (danjs):

put both fractions over a common denominator, 15

OpenStudy (minthia):

1/5 and 1/3?

OpenStudy (danjs):

\[\frac{ -1 }{ 3 }*\frac{ 5 }{ 5 }= \frac{ -5 }{ 15 }\]

OpenStudy (danjs):

\[\frac{ -2 }{ 5 }*\frac{ 3 }{ 3 } = \frac{ -6 }{ 15 }\]

OpenStudy (danjs):

since they are both divided by the same denominator, you can just compare the numerators with > < =

OpenStudy (danjs):

It is better for more strange fractions that arent common

OpenStudy (minthia):

So the answer would be <

OpenStudy (danjs):

no, the more negative , the less it is

OpenStudy (minthia):

Oh so >

OpenStudy (danjs):

\[\frac{ -5 }{ 15 }>\frac{ -6 }{ 15 }\]

OpenStudy (minthia):

thank you

OpenStudy (danjs):

welcome